# WinkTPG: An Execution Framework for Multi-Agent Path Finding Using Temporal Reasoning

### v1

@@ 2.2 Temporal Plan Graph (TPG)
2.3 Related Works
MAPF methods have been developed to find collision-free paths for a large group of agents. Leading methods, such as EECBS (Li, Ruml, and Koenig 2021), PIBT (Okumura et al. 2022), and PBS (Ma et al. 2019), can find collision-free paths for thousands of agents. However, these methods assume all agents move synchronously, which is unrealistic in real-world scenarios due to their reliance on simplified kinodynamic models. To address this issue, some methods (Andreychuk et al. 2021; Yan and Li 2024; Moldagalieva et al. 2024) integrate more accurate kinodynamic models during planning. Nevertheless, these methods often face scalability challenges due to the increased computational complexity introduced by the kinodynamic models. Another group of methods (Hönig et al. 2016; Zhang et al. 2021) first generates a MAPF plan with a simplified kinodynamic model and subsequently derives speed profiles with the accurate model based on these paths. For instance, KDN (Zhang et al. 2021) employs a TPG to encode the passing order constraints defined in the MAPF plan. It then uses a Mixed-Integer Linear Programming (MILP) solver to find kinodynamically feasible speed profiles for agents based on the TPG.
To handle time uncertainty that exists during execution, some methods (Atzmon et al. 2020; Peltzer et al. 2020) account for uncertainty during planning. These methods first build a model of the uncertainty and then plan collision-free paths, ensuring that agents remain collision-free even when time uncertainties occur as described by the models. While effective in some cases, these methods often lead to poor performance due to being overly conservative during the planning stage, as they cannot access the actual delays that occur during execution. Additionally, these methods are often limited in scalability.
Introducing an execution framework to MAPF is a promising approach to meet robot kinodynamic constraints while ensuring robustness. Action Dependency Graph (ADG) (Hönig et al. 2019) uses a TPG to ensure the passing order and defines an execution framework for speed profile generation. Specifically, it categorizes the vertices in the graph into three statuses: staged, enqueued, and finished. A vertex is marked as staged if the agent cannot move to it yet. It transitions to enqueued when the agent can safely move here. This happens only if it has no preceding vertices or if the following conditions are met: (1) its preceding vertex connected by a Type-1 edge is enqueued or finished, and (2) all preceding vertices connected by Type-2 edges are finished. Once a vertex transitions to enqueued, ADG generates a speed profile for the agent to reach the vertex. A vertex is marked as finished when the agent has reached it. ADG dynamically updates vertex statuses during execution and provides robustness guarantees under temporal uncertainty.
However, ADG only plans the speed profile for a vertex after it transitions to enqueued, even when it is predictable that this transition will happen before the agent reaches it. This conservative strategy can lead to unnecessary deceleration. As shown in Fig. 2, R2 (blue) must wait until R1 (green) leaves cell B2 before it can proceed. Although it is evident early on that R1 will leave B2 before R2 arrives, ADG does not allow R2 to commit to the optimal speed profile until the dependency is officially satisfied. This results in R2 decelerating prematurely (top chart in (c)), compared to the optimal case (bottom chart) where it could have maintained maximum speed. The problem is worsened by potential communication delays in real-world settings, where agents rely on frequent updates from a central server. Consequently, this overly conservative strategy can result in poor performance in many scenarios.
3 Kinodynamic TPG (kTPG)
Definition 1 below specifies the problem addressed in this section. In Sections 4 and 5, we will extend this definition to account for temporal uncertainty during agent movement.
Definition 1 (Multi-Agent Execution Problem (MAEP))
Given a MAPF plan, the objective of MAEP (Zhang et al. 2021) is to generate a set of speed profiles for all agents to follow during execution. These profiles must satisfy the kinodynamic constraints imposed by the agents’ physical models, remain collision-free, and respect the precedence relationships defined in the MAPF plan. The objective is to minimize the total time for all agents to reach their goals.
3.1 kTPG Overview
The precedence constraints in a MAPF plan require that each agent (1) follows its assigned path in sequence and (2) respects the relative passing ordering of shared locations with other agents. These constraints can be encoded in a TPG $\mathcal{G}$. Assume each agent $a_{i}$ begins at its start location $q_{i}^{0}=q_{i}^{s}$ at time $t=0$, and its state at time $t$ is defined as $s_{i}(t)=(q_{i}^{t},\theta_{i}^{t})$, with $q_{i}^{t}\in\mathcal{V}_{U}$ and $\theta_{i}^{t}\in\Theta$ being its location and orientation, respectively. We use reach time $t_{i}^{r}(q_{i}^{k})$ to denote the time when $a_{i}$ reaches location $q_{i}^{k}$. For ease of reading, we introduce the term leave time $t^{l}(q_{i}^{k})=t^{r}(q_{i}^{k+1})$, i.e., the time when $a_{i}$ reaches its next location. We say $a_{i}$ occupies location $q_{i}^{k}$ during $[t_{i}^{r}(q_{i}^{k}),t_{i}^{r}(q_{i}^{k+1})]$. A collision occurs if two agents occupy the same location during overlapping time intervals. We say that the speed profiles satisfy the edges in $\mathcal{G}$ if the time intervals during which agents occupy each location (as defined by the speed profiles) do not violate the precedence constraints imposed by the edges. It is easy to show that satisfying all edges in the TPG guarantees that the precedence constraints defined in the MAPF plan are preserved. The high-level idea of kTPG is to introduce a reserved interval constraint on each vertex of the TPG, initialized as [0,)[0$\infty$)[ 0 , ∞ ). It then generates speed profiles for each agent independently that visit each vertex within its corresponding reserved intervals. kTPG iteratively updates these constraints and the corresponding speed profiles until the reserved intervals of the vertices corresponding to the same location do not overlap, which in turn ensures that the corresponding speed profiles do not lead to collisions.
We use the example in Fig. 3 to show a single iteration of this process. Each iteration begins by selecting an agent to replan, following the selection rule in Section 3.2; in this case, $a_{2}$ is selected. We define a Type-2 edge as satisfied if the reserved intervals at its associated vertices do not overlap and the visiting order is preserved. Otherwise, it is conflicting. All Type-2 edges in our example are conflicting in this iteration. We then generate a speed profile for $a_{2}$ from its first to last unlocked vertex, i.e., vertices without any incoming conflicting edges, using the method in Section 3.3. These vertices are shown as dark green circles in Fig. 3 (b). Next, we identify the conflicting edges originating from these unlocked vertices (i.e., the red edge in Fig. 3 (b)). We use the generated speed profile to derive a split time for the target vertex (here, C2) and update the reserved intervals for both agents accordingly so that the intervals no longer overlap (see Section 3.4). This update allows us to satisfy a conflicting edge and unlock additional vertices of other agents, as shown in Fig. 3 (c). In the next iteration, we use the newly unlocked vertices to satisfy more Type-2 edges by selecting another agent and updating its speed profile. This process repeats until all Type-2 edges are satisfied, at which point we get complete, collision-free speed profiles for all agents.
@@ 3.2 Selecting Agent
3.3 Updating Speed Profile
@@
After selecting an agent, we compute a speed profile that visits each unlocked vertex within its reserved interval, starts and ends with zero speed, and satisfies the robot’s kinodynamic constraints. Because these unlocked vertices have no incoming conflicting edges, the resulting speed profile is guaranteed to be collision-free and can safely be used to update the reserved intervals. This can be accomplished using a single-agent motion planner or speed optimization techniques (Liu, Zhan, and Tomizuka 2017; Ali and Yakovlev 2023). Empirically, we achieve this by using a 1-D adaptation of SIPP-IP (Ali and Yakovlev 2023), an optimal algorithm that performs an A*-type search along the vertices using a set of predefined motion primitives.
3.4 Satisfying Conflicting Edges
@@
After we get the speed profile for $a_{i}$, we update the reserved intervals to satisfy conflicting edges. Specifically, for each conflicting edge $(v_{i}^{k+1},v_{j}^{s})$ originating from an unlocked vertex $v_{i}^{k+1}$ of $a_{i}$, we use the leave time of $a_{i}$ at the shared location $loc(v_{i}^{k})=loc(v_{j}^{s})$, denoted as $t^{l}$, as the split time to divide the time interval at the shared location into $[0,t^{l}]$ and $(t^{l},\infty)$. We enforce that $a_{i}$ and $a_{j}$ occupy the shared location within these intervals respectively. Concretely, we update the reserved intervals for $v_{i}^{k}$ and $v_{j}^{s}$ by intersecting them with the corresponding segments.
@@
Fig. 3 shows an example. Once the speed profile for agent $a_{2}$ is obtained, we find all conflicting edges originating from unlocked vertices of $a_{2}$, which is the first Type-2 edge from $a_{2}$ to $a_{1}$, with shared location $C_{2}$. From the speed profile of $a_{2}$, we use the leave time $t_{2}^{l}(C_{2})$ as the split time. We divide the time interval at $C_{2}$ into $[0,t_{2}^{l}(C_{2})]$ and $(t_{2}^{l}(C_{2}),\infty)$. The reserved intervals of $a_{1}$ and $a_{2}$ are updated by intersecting their current reserved intervals with these segments, respectively. By assigning these intervals, the Type-2 edge is marked as satisfied. Any speed profile generated for $a_{2}$ in subsequent iterations must adhere to this updated interval.
@@
4 kTPG with Uncertainty
In the previous section, we assumed the agents can perfectly follow the speed profiles. However, in realistic scenarios, agents may experience temporal uncertainty during movement due to slippage or controller variability. To address this, we first introduce an uncertainty model to represent these temporal deviations. We then present kTPGu (kTPG with Uncertainty), which extends kTPG to account for these uncertainties using a safety margin mechanism.
Definition 2 (MAEP under Uncertainty)
We extend the MAEP in Definition 1 by incorporating an uncertainty model for movement. The expected time for $a_{i}$ to move from $q_{i}^{k}$ to $q_{i}^{k+1}$ is denoted $T_{i}^{k}$. However, due to the temporal uncertainty, the actual move time includes a random noise modeled as a normal distribution with zero mean and variance proportional to the movement distance. Thus, the actual move time $\hat{T}_{i}^{k}$ from $q_{i}^{k}$ to $q_{i}^{k+1}$ is given by:
$\hat{T}_{i}^{k}=T_{i}^{k}+\mathcal{N}(0,(\varepsilon_{i}^{k})^{2}),\;({\varepsilon_{i}^{k}})^{2}=K_{i}\cdot d_{k}^{k+1},$ (1)
where $d_{k}^{k+1}$ is the distance between $q_{i}^{k}$ and $q_{i}^{k+1}$, and $K_{i}$ is an agent-specific hyper-parameter.
To prevent collisions resulting from temporal uncertainty, we introduce safety margins to extend kTPG. Intuitively, a safety margin is an extra time buffer added between reserved intervals at shared locations to account for early or delayed arrivals, ensuring agents remain collision-free.
Using our uncertainty model, we can represent the time $a_{i}$ reaches a certain vertex $v_{i}^{k}$ as a normal distribution:
$t_{i}^{r}(v_{i}^{k})=t_{i}^{r}(v_{i}^{0})+\sum_{j=0}^{k-1}\hat{T}_{i}^{j}=\mathcal{N}(\mu_{i}^{k},(\sigma_{i}^{k})^{2}).$ (2)
with $\mu_{i}^{k}=\mu_{i}^{0}+\sum_{j=0}^{k-1}T_{i}^{j}$, $(\sigma_{i}^{k})^{2}=(\sigma_{i}^{0})^{2}+\sum_{j=0}^{k-1}(\varepsilon_{i}^{j})^{2}$. The reach time at the first vertex $v_{i}^{0}$ is $\mathcal{N}(0,0)$, as agents start precisely at $t=0$. For a Type-2 edge $(v_{i}^{m+1},v_{j}^{n})$ between agents $a_{i}$ and $a_{j}$ with $loc(v_{i}^{m})=loc(v_{j}^{n})$, the probability of maintaining its passing order is:
$P(t_{i}^{l}(v_{i}^{m})<t_{j}^{r}(v_{j}^{n}))$ $=P(t_{i}^{r}(v_{i}^{m+1})-t_{j}^{r}(v_{j}^{n})<0)$ (3)
$=P(\mathcal{N}(0,(\sigma_{i}^{m+1})^{2}+(\sigma_{j}^{n})^{2})<\Delta\mu)$
$=F_{\mathcal{N}_{m+1}^{n}}(\Delta\mu).$
where $\Delta\mu=\mu_{j}^{n}-\mu_{i}^{m+1}$, $F_{\mathcal{N}_{m+1}^{n}}$ is the cumulative density function of $\mathcal{N}_{m+1}^{n}=\mathcal{N}(0,(\sigma_{i}^{m+1})^{2}+(\sigma_{j}^{n})^{2})$. Since $\sigma$ depends only on the travel distance, we can derive $F_{\mathcal{N}_{m+1}^{n}}$ given the paths of the agents. To ensure with probability at least $P_{d}$ that the passing order is not violated, i.e., $P(t_{i}^{l}(v_{i}^{m})<t_{j}^{r}(v_{j}^{n}))\geq P_{d}$, we can derive:
$\mu_{j}^{n}-\mu_{i}^{m+1}$ $\geq F_{\mathcal{N}_{m+1}^{n}}^{-1}(P_{d}).$ (4)
From Eq. 4, the passing order defined by a Type-2 edge is maintained with a probability greater than $P_{d}$ if the expected leave time of $a_{i}$ exceeds the expected reach time of $a_{j}$ by at least $F_{\mathcal{N}_{m+1}^{n}}^{-1}(P_{d})$. Thus, we introduce kTPGu to handle time uncertainty by adding a minimum time gap between the agents’ expected leave and reach times. Compared to the kTPG, the only modification is in updating reserved intervals. Once the leave time $t_{i}^{l}(v_{i}^{m})=t_{i}^{r}(v_{i}^{m+1})$ for $a_{i}$ at $v_{i}^{m}$ is determined, we use Eq. 3 to compute the earliest time when $a_{j}$ can reach $v_{j}^{n}$. The difference between $t_{i}^{l}(v_{m})$ and $t_{j}^{r}(v_{j}^{n})$ is referred to as the safety margin. Using this safety margin, we divide the time at $loc(v_{i}^{m})$ into two intervals: $[0,t_{i}^{r}(v_{i}^{m+1})]$ for $a_{i}$ and $(t_{i}^{r}(v_{i}^{m+1})+F_{\mathcal{N}_{m+1}^{n}}^{-1}(P_{d}),\infty)$ for $a_{j}$. These intervals are then intersected with the existing reserved intervals of $a_{i}$ and $a_{j}$, resulting in updated reserved intervals for both agents, ensuring the passing order is maintained under time uncertainty.

### v2

@@ III-C Selecting Agent
Input: MAPF plan $P=\{p_{i}\}_{i=1}^{I}$, kinodynamic models $\{\mathcal{X}_{i}\}$
Output: Speed profiles $\{x_{i}(\cdot)\}_{i=1}^{I}$
1
2Construct TPG $\mathcal{G}$ from $P$ and initialize reserved intervals $\mathcal{R}$;
Initialize $\mathcal{C}\leftarrow\emptyset$ ;
// $\mathcal{C}$: agents with complete speed profiles to their goals
3 Mark all Type-2 edges in $\mathcal{G}$ as unsatisfied;
4while $|\mathcal{C}|<I$ do
5   Select an agent $a_{i}$ to refine;
6   Determine the set of unlocked vertices on $a_{i}$’s path;
7   Compute a kinodynamically feasible speed profile $x_{i}(\cdot)$ on the unlocked vertices, subject to $\mathcal{X}_{i}$ and $\mathcal{R}$;
8   Update the reserved intervals $\mathcal{R}$;
9   Mark any Type-2 edges that have become satisfied as satisfied;
10 if $x_{i}(\cdot)$ is computed up to the goal then
11    $\mathcal{C}\leftarrow\mathcal{C}\cup\{i\}$;
12
13
return $\{x_{i}(\cdot)\}_{i=1}^{I}$
Algorithm 1 kTPG
III-D Updating Speed Profile
@@
After selecting an agent, we compute a speed profile for its 1-D path that (i) visits each unlocked vertex within its reserved interval, (ii) starts and ends with zero speed, and (iii) satisfies kinodynamic constraints. Because the unlocked vertices have no incoming conflicting edges, any speed profile that respects the reserved intervals is collision-free with respect to already-fixed agents, and can be used to update the reserved intervals.
We formulate this step as a single-agent planning problem in time–distance space along the fixed geometric path: each vertex induces a safe time window in which the agent may occupy the corresponding shared location, and each edge traversal must respect the agent’s kinodynamic limits. This problem can be addressed using MILP-based trajectory optimization methods [16, 17] or search-based motion primitives methods [11, 1]. Empirically, we achieve this by using a 1-D adaptation of SIPP-IP [1], an optimal algorithm that performs an A*-type search along the vertices using a set of predefined motion primitives. Given the predefined motion primitives, SIPP-IP is complete and optimal, i.e., it provably finds the speed profile that minimizes the reach time at the last unlocked vertex if one exists and returns failure otherwise. Importantly, the proposed framework is not tied to a particular speed-profile generation method. It requires only a single-agent planner capable of producing a speed profile that respects the reserved time intervals and satisfies the kinodynamic constraints encoded in $\mathcal{X}_{i}$. Alternative kinodynamic models, such as jerk-limited $S$-curve profiles, can therefore be incorporated by modifying the underlying primitive set or replacing the trajectory optimizer, without altering the temporal reasoning mechanism or the reserved-interval update rules of kTPG.
III-E Satisfying Conflicting Edges
@@
After obtaining the speed profile for $a_{i}$, we update the reserved intervals to satisfy conflicting edges. Specifically, for each conflicting edge $(v_{i}^{k+1},v_{j}^{s})$ originating from an unlocked vertex $v_{i}^{k+1}$ of $a_{i}$, we use the leave time of $a_{i}$ at the shared location $loc(v_{i}^{k})=loc(v_{j}^{s})$, denoted as $t^{l}$, as the split time to divide the time interval at the shared location into $0,t^{l}]$ and $(t^{l},\infty)$. We enforce that $a_{i}$ and $a_{j}$ occupy the shared location within these intervals respectively. Concretely, we update the reserved intervals for $v_{i}^{k}$ and $v_{j}^{s}$ by intersecting them with the corresponding segments.
@@
Fig. 4 shows an example. Once the speed profile for agent $a_{2}$ is obtained, we find all conflicting edges originating from unlocked vertices of $a_{2}$, which is the first Type-2 edge from $a_{2}$ to $a_{1}$, with shared location $C_{2}$. From the speed profile of $a_{2}$, we use the leave time $t_{2}^{l}(C_{2})$ as the split time. We divide the time interval at $C_{2}$ into $0,t_{2}^{l}(C_{2})]$ and $(t_{2}^{l}(C_{2}),\infty)$. The reserved intervals of $a_{1}$ and $a_{2}$ are updated by intersecting their current reserved intervals with these segments, respectively. By assigning these intervals, the Type-2 edge is marked as satisfied. Any speed profile generated for $a_{2}$ in subsequent iterations must adhere to this updated interval.
@@
IV kTPG with Uncertainty
In the previous section, execution was assumed deterministic. In realistic deployments, agents experience (execution timing) uncertainty due to controller noise, actuation latency, slippage, and other execution-level perturbations. To account for such temporal uncertainty, we extend the Multi-Agent Execution Problem (MAEP) to include uncertainty models and derive model-conditional safety margins that ensure deterministic or probabilistic precedence satisfaction under those models. These safety margins are integrated into kTPG to form kTPGu, which maintains collision-free execution under the adopted uncertainty regimes.
IV-A MAEP under Uncertainty
We extend the MAEP in Section III-A by incorporating an execution timing uncertainty model for movement. Concretely, we model such uncertainty as additive deviations to nominal traversal times. Specifically, if $T_{i}^{k}$ is the nominal time for agent $a_{i}$ to move from $q_{i}^{k}$ to $q_{i}^{k+1}$, then the realized time is
$\hat{T}_{i}^{k}=T_{i}^{k}+\delta_{i}^{k},$ (1)
where $\delta_{i}^{k}$ is a temporal perturbation.
In this work, we adopt three representative execution timing uncertainty models from the MAPF, stochastic control, and temporal reasoning literature:
  1. 1.
    Bounded Cumulative Delay Model. This model represents bursty but bounded execution delays, such as those caused by transient control saturation or local disturbances. Bounded cumulative delay models have been used in MAPF under uncertainty because they impose a finite delay budget per agent without requiring full stochastic assumptions [3]. Formally,
    $\delta_{i}^{k}=\begin{cases}\eta&\text{with prob. }p,\ \text{if }\sum_{j=0}^{k-1}\delta_{i}^{j}<\mathcal{K}_{i},\\ 0&\text{otherwise},\end{cases}$
    where $p$ is the probability of delay per-move, $\eta>0$ is the magnitude of the unit delay, and $\mathcal{K}_{i}$ is the total accumulated delay limit for the agent. Intuitively, each time an agent moves, it may experience a fixed delay $\eta$ with probability $p$. However, the total accumulated delay across all moves is capped at $\mathcal{K}_{i}$. We assume $\mathcal{K}_{i}$ is a multiple of $\eta$, i.e., $\mathcal{K}_{i}=\ell_{i}\cdot\eta$ for some $\ell_{i}\in\mathbb{N}$.
  2. 2.
    Bounded Uniform Noise Model. In temporal reasoning under bounded intervals (e.g., Simple Temporal Networks), traversal times are assumed to lie within a known interval [7]. We model this as
    $\delta_{i}^{k}\sim\mathcal{U}(-\epsilon_{i},\epsilon_{i}),$
    where $\mathcal{U}(-\epsilon_{i},\epsilon_{i})$ denotes the uniform distribution over $\epsilon_{i},\epsilon_{i}]$, and $\epsilon_{i}>0$ defines the maximum deviation from nominal traversal time.
  3. 3.
    Gaussian Noise Model. We model aggregate timing variability arising from many independent small perturbations as additive zero-mean Gaussian noise. Gaussian noise models are widely used in stochastic control and robotics to represent unbounded execution uncertainty [27]:
    $\delta_{i}^{k}\sim\mathcal{N}(0,\sigma^{2}),$
    where $\mathcal{N}(0,\sigma^{2})$ denotes the normal distribution with zero mean and variance $\sigma^{2}$, and the variance scales with the traversal distance.
Unless otherwise stated, perturbations $\delta_{i}^{k}$ are assumed independent across agents and time. Correlated and heavy-tailed noise are outside the scope of the current derivation.
IV-B Probabilistic Safety Margin Formulation
Let $\mu_{i}^{k}=\sum_{j=0}^{k-1}T_{i}^{j}$ denote the nominal reach time of $a_{i}$ at vertex $v_{i}^{k}$, and let $\Delta_{i}^{k}=\sum_{j=0}^{k-1}\delta_{i}^{j}$ be the cumulative perturbation. Then the actual reach time is
$t_{i}^{r}(v_{i}^{k})=\mu_{i}^{k}+\Delta_{i}^{k}.$ (2)
For a Type-2 edge $(v_{i}^{m+1},v_{j}^{n})$ between agents $a_{i}$ and $a_{j}$ with $loc(v_{i}^{m})=loc(v_{j}^{n})$, we require that the precedence constraint is satisfied with probability at least $P_{d}$, the safety probability threshold:
$P\bigl(t_{i}^{l}(v_{i}^{m})<t_{j}^{r}(v_{j}^{n})\bigr)\;\geq\;P_{d}.$ (3)
Under the adopted uncertainty model and independence assumptions, this probabilistic constraint can be converted into a safety margin $\gamma_{i,j}^{m,n}(P_{d})$ such that
$\mu_{j}^{n}-\mu_{i}^{m+1}\;\geq\;\gamma_{i,j}^{m,n}(P_{d}).$ (4)
The safety margin depends on the distribution (or bounds) of the difference $\Delta_{i}^{m+1}-\Delta_{j}^{n}$ under the selected uncertainty model. Since the first two models have bounded perturbations, we can derive safety margins that guarantee deterministic precedence satisfaction, i.e., $P_{d}=1$. The Gaussian model has unbounded perturbations and thus only admits probabilistic guarantees for a given $P_{d}<1$.
  1. 1.
    Bounded Cumulative Delay Model. Under the bounded cumulative delay model, each agent’s delay at each step is either 0 or $\eta$, and the total delay budget is $\mathcal{K}_{i}$ for agent $a_{i}$ and $\mathcal{K}_{j}$ for agent $a_{j}$. Therefore, the cumulative perturbations satisfy
    $\Delta_{i}^{m+1}-\Delta_{j}^{n}\;\in\;[-\min(n,\ell_{j})\,\eta,\ \min(m{+}1,\ell_{i})\,\eta].$ (5)
    Thus, a valid deterministic safety margin ($P_{d}=1$) is:
    $\gamma_{i,j}^{m,n}(1)=\min(m{+}1,\ell_{i})\,\eta,$ (6)
    which accounts for the maximum possible prefix delay agent $a_{i}$ could incur under its finite budget.
  2. 2.
    Bounded Uniform Noise Model. When $\delta_{i}^{k}\sim\mathcal{U}(-\epsilon_{i},\epsilon_{i})$ and $\delta_{j}^{k}\sim\mathcal{U}(-\epsilon_{j},\epsilon_{j})$, each term $\delta_{i}^{k}$ and $\delta_{j}^{k}$ lies within $\epsilon,\epsilon]$. Consequently,
    $\Delta_{i}^{m+1}-\Delta_{j}^{n}\;\in\;[-(m{+}1)\epsilon_{i}-n\epsilon_{j},\,(m{+}1)\epsilon_{i}+n\epsilon_{j}].$ (7)
    For deterministic precedence satisfaction ($P_{d}=1$), a valid safety margin is
    $\gamma_{i,j}^{m,n}(1)=(m{+}1)\epsilon_{i}+n\epsilon_{j},$ (8)
    which corresponds to the worst-case bound.
  3. 3.
    Gaussian Noise Model. Under additive zero-mean Gaussian noise, if $\delta_{i}^{k}\sim\mathcal{N}(0,\sigma_{i}^{2})$ and $\delta_{j}^{k}\sim\mathcal{N}(0,\sigma_{j}^{2})$ independently, then $\Delta_{i}^{m+1}\sim\mathcal{N}(0,(m{+}1)\sigma_{i}^{2})$ and $\Delta_{j}^{n}\sim\mathcal{N}(0,n\sigma_{j}^{2})$. The difference $\Delta_{i}^{m+1}-\Delta_{j}^{n}$ is therefore Gaussian with variance
    $\mathrm{Var}[\Delta_{i}^{m+1}-\Delta_{j}^{n}]=(m{+}1)\sigma_{i}^{2}+n\sigma_{j}^{2}.$ (9)
    The $(1-P_{d})$-quantile of this distribution gives
    $\gamma_{i,j}^{m,n}(P_{d})=\Phi^{-1}(P_{d})\cdot\sqrt{(m{+}1)\sigma_{i}^{2}+n\sigma_{j}^{2}},$ (10)
    where $\Phi^{-1}$ is the inverse cumulative distribution function of the standard normal.
IV-C Integration with kTPG
To integrate uncertainty into the planning framework, kTPGu enforces a minimum time gap between agents’ expected leave and reach times at shared locations based on the computed safety margins. Specifically, once $t_{i}^{l}(v_{i}^{m})=t_{i}^{r}(v_{i}^{m+1})$ is fixed for agent $a_{i}$, the earliest permissible reach time for agent $a_{j}$ at $v_{j}^{n}$ is
$t_{j}^{r}(v_{j}^{n})\;\geq\;t_{i}^{r}(v_{i}^{m+1})+\gamma_{i,j}^{m,n}(P_{d}).$ (11)
Reserved intervals are then updated by intersecting these constrained windows with the existing feasibility intervals for both agents, thereby maintaining the precedence under temporal uncertainty. This mechanism preserves collision-free execution under the adopted uncertainty models.
Remark. The choice of execution timing uncertainty model depends on the robot platform, controller, and operating environment. Our framework is mainly intended to handle execution delays arising from routine variability, such as tracking errors or local disturbances, which are typically bounded or light-tailed. More severe heavy-tailed disturbances, e.g., robot breakdowns or major system faults, are not explicitly modeled here and should instead be handled by higher-level path replanning or task reassignment.